OpenStudy (anonymous):

What is the range of ƒ(x)=3x-5 when the domain is {-6, 4, 8}?

OpenStudy (cwrw238):

just plug in x = -6, x=4 and x = 8 into f(x) and work out the 3 results

OpenStudy (anonymous):

So, f(-6)=3(-6)-5 f(4)=3(4)-5 f(8)=3(8)-5 -23, 7 and 19?

OpenStudy (cwrw238):

yes - the range is written {-23,7,19}

OpenStudy (anonymous):

Thank you so much!! xxx

OpenStudy (cwrw238):

yw